OpenStudy (anonymous):

What is the associative property of 1/6(6x)? Will fan and medal for the help.

OpenStudy (misty1212):

HI!!!

OpenStudy (misty1212):

is it \[\frac{1}{6}(6x)\]

OpenStudy (anonymous):

Yes sorry it's been a long day lol I just got out of a chemistry class and my sister asked for math help and I just brain farted. Thanks so much!

OpenStudy (misty1212):

in the associative law,the symbols stay in the same order, but the parentheses move i.e \[\left(\frac{1}{6}\times 6\right)x\]

OpenStudy (misty1212):

i guess they are trying to train her to say \[\frac{1}{6}\times (6x)=x\]